# Aligning Individual and Collective Objectives in Multi-Agent Cooperation

### v1

@@ 5.1 Two-Player Public Goods Game
A two-player Public Goods matrix game ${\mathcal{G}}$ is widely utilized to study cooperation in social dilemmas. The game involves players $\{1,2\}$ with parameters $\{{\bm{w}}_{1},{\bm{w}}_{2}\}$ and payoffs $\{p_{1},p_{2}\}$. The social welfare, denoted as $SW=p_{1}+p_{2}$. Each player, i.e., $i\in\{1,2\}$, contributes an amount $a_{i}$ within a budgeted range $[0,b]$, and the host evenly distributes these contributions as $\frac{c}{2}(a_{1}+a_{2})$, where $1<c\leq 2$. Consequently, each player’s payoff $p_{i}(a_{1},a_{2})$ is estimated as $b-a_{i}+\frac{c}{2}(a_{1}+a_{2})$. In our experiments, we set the budget $b$ to $1$ and weight $c$ to $1.5$, with the Nash equilibrium of game at $(0,0)$.
Results. Table 1 compares averindividual rewards, social welfare, and the equality metric sourced from 50 random runs, each capped at 100 steps. The rows $r_{1},r_{2}$ and $SW$ represent the rewards accrued by each individual player and the collective group respectively. Row $E$ denotes the equality metric. The table reveals that AgA surpasses benchmark methods in individual rewards and social welfare, and equality metric. The social optimum of the game is at coordinates $(1,1)$, which corresponds to a maximum achievable social welfare of $3$. Among all the algorithms, AgA achieves the closest social welfare score to this optimum, with a value of 2.90. In addition, AgA demonstrates the highest level of equality, with a score of 0.96, when compared to the baseline algorithms. Additionally, Fig. 6 in Appendix B visualizes the action distributions of these methods, demonstrating convergence points, further supporting that AgA converges more efficiently towards the social optimum.
@@ 5.2 Sequential Social Dilemma: Cleanup and Harvest
We conduct experiments in sequential social dilemma environments with a group size of five namely, Harvest and Cleanup (Hughes et al., 2018). In Cleanup, the agents primarily derive rewards from harvesting apples in a nearby orchard, which is contingent upon river pollution levels. If negligence towards the escalating pollution levels surpasses a threshold, apple growth ceases, hence leading to a conflict of interest between obtaining individual rewards and safeguarding joint welfare. Harvest’s premise, on the other hand, involves agents receiving rewards by gathering apples, whereby optimal apple regrowth is dependent on the presence of adjacent apples. A collective dilemma appears when rampant harvesting reduces the apple regeneration rate, thus decrementing overall team rewards. Simul-Ind, Simul-Co, CGA, and SVO are benchmarked against our AgA methodology in our experiments. The implementation of the Simul-Ind algorithm is based on the IPPO algorithm (de Witt et al., 2020), with all other strategies leveraging the shared parameters of the IPPO algorithm as well. The joint reward formula used for the calculation of collective loss for Simul-Co, CGA, and AgAs is designed to promote both improved collective performance and equitable behavior among agents, i.e, $\sum_{i}\left(r_{i}-\alpha(1-actan\left(\frac{sum_{j,j\neq i}r_{j}}{r_{i}}\right)\right)$ with constant $\alpha$. More detailed information on these environments and algorithm execution can be found in Appendix C.1.
Envs Metrics Algorithms
Simul-Ind Simul-Co SVO CGA AgA
Cleanup $SW$ 24.86 72.02 27.30 44.23 83.90
$E$ 0.97 0.88 0.85 0.91 0.95
Harvest $SW$ 623.13 625.80 638.95 667.37 683.42
$E$ 0.97 0.96 0.97 0.98 0.98
Table 2: The highest social welfare achieved throughout training along with the corresponding equality metric. The foremost row for each environment reflects the best social welfare, while the second row indicates the associated equality metric.
Results. Fig. 3(a) and Fig. 3(c) present the average episodic social welfare comparison of the AgA algorithm and AgA/Sign, excluding the sign selection (Corollary 4.3). These figures also highlight the effect of varying the alignment parameter $\lambda$. The prescribed $\lambda$ values for AgA/Sign is 100 for Cleanup and 0.1 for Harvest, resulting in the best performance for each. Failing to select signs in AgA/Sign leads to poorer collective results. Fig. 3(b) and Fig. 3(d) compare the social welfare achieved by AgA, with the optimal $\lambda$ value, with baseline methods. Here, AgA $(\lambda=100)$ in Cleanup and AgA $(\lambda=0.1)$ in Harvest surpass the benchmark methods for the majority of the training process. Table 2 presents the peak social welfare achieved during training alongside the equality metric. For simplicity, the AgA column denotes the respective optimal $\lambda$ for each scenario. In Cleanup, AgA surpasses the second-best method, Simul-Co, by over 10 points. However, although Simul-Ind shows superior equality performance, it delivers the lowest social welfare. Contrarily, AgA achieves a satisfactory balance between equity and team performance. Analogous conclusions apply to the Harvest, where AgA outperforms baselines in social welfare and equality.
(a) MMM2: AgA with various $\lambda$ and AgA/Sign
(b) MMM2: comparison between AgA and baselines
Figure 5: The comparison of win rates on the mixed-motive MMM2 map of SMAC. The left figure shows results of varying alignment parameter $\lambda$ values in AgA, as well as AgA sans sign selection (AgA/Sign), The right figure compare the win rate of AgA $(\lambda=100)$ with baselines. The bold lines illustrate the mean of the win rate over 3 seeds, while the encompassing shaded regions depict the standard deviation.
5.3 Mixed-motive SMAC
Despite their widespread use in the study of mixed-motive games, sequential social dilemma games can be relatively simplistic compared to other testbeds in the field of MARL, particularly when considering real-world scenarios. Cooperative Multi-Agent Reinforcement Learning often adopts more intricate environments such as SMAC (Samvelyan et al., 2019). In this game environment, agents must collaborate to defeat their opponents. We modify the MMM2 map from SMAC as a mixed-motive testbed. Here, agents command various game entities confronting opposing forces. Unlike the original StarCraft II game’s team-oriented architecture, we instill individual interests in each agent to enhance mixed-motive characteristics. Reward mechanisms have been modified such that an agent’s individual reward aligns with the damage it inflicts on the enemy, replacing the standard practice of distributing total damage across all agents. Penalties for agent removal also reinforce self-preservation, accentuating inherent agent selfishness. The experimental implementation utilizes the IPPO algorithm (de Witt et al., 2020) for Simul-Ind, while employing the MAPPO framework (Yu et al., 2022) for Simul-Co, SVO, CGA, and AgA. More details can be found in Appendix C.2.
Results. Fig. 4(a) compares average win rates for various $\lambda$ parameters in AgA, also encompassing AgA without sign selection, as stated in Corollary 4.3. The results suggest comparable performance for $\lambda=1$ and $100$, with a lower value of $0.1$ trailing noticeably. Notably, AgA with $\lambda=0.1$ underperforms, even when contrasted with AgA’s ablation study. Fig. 4(b) benchmarks the win rate for AgA $(\lambda=100)$ against Simul-Ind, Simul-Co, CGA, and SVO. Within this context, Simul-Ind, implemented via IPPO, struggles with the mixed-motive MMM2 task. Impressively, our AgA $(\lambda=100)$ method achieves a win rate peak of $0.85$ (standard deviation of $0.09$), significantly outperforming the SVO method that only nets $44.55$.
@@ 6 Conclusion
In this paper, we propose AgA, an gradient adjustment method, specifically designed to align individual and collective objectives within mixed-motive cooperation scenarios. To materialize this, we first model the mixed-motive game as a differentiable game, thereby facilitating the analysis of learning dynamics. Thereafter, we analyze the optimization paths in a toy differentiable mixed-motive game. We reveal that optimization techniques that heavily rely on collective loss suffer from inherent limitations. These approaches often overlook individual preferences, and existing gradient adjustment methods are inadequate for addressing mixed-motive cooperation challenges. Our AgA strategy efficiently bridges the gap between individual and collective objectives. The optimization paths further support our theoretical claims by demonstrating that careful selection of the sign of the alignment weight enables convergence towards stable fixed points while escaping unstable ones. Comprehensive experiments conducted on mixed-motive testbeds and modified SMAC environments provide further evidence of the superior performance of AgA. It consistently outperforms existing baselines across multiple evaluation metrics.
Limitations and Future Work. The experiments conducted in this study were confined to a single map of the SMAC environment. Our future objectives include investigating the mixed-motive cooperative mechanism in real-world scenarios and developing a more complex mixed-motive testbed. In this paper, while we strive to align the objectives of both individuals and the collective, our primary objective is to converge towards a stable equilibrium of the collective objective. For future research, we intend to broaden our comprehension of how individual objectives interact in mixed-motive games, placing a greater emphasis on understanding their dynamics in mixed-motive cooperation.
Impact Statements
@@
This paper presents a novel approach in the field of multi-agent learning, aimed at advancing collaborations with mixed motivations. The proposed framework addresses a fundamental challenge in machine learning, wherein agents with diverse objectives are required to cooperate to achieve collective goals. The work described in this paper holds significant potential for various societal applications. By enabling effective cooperation among agents with different motivations, our approach can be leveraged in domains such as games, autonomous systems, smart cities, and decentralized resource allocation. Such advancements have numerous societal consequences, including improved efficiency in transportation systems, optimized resource allocation in energy grids, and enhanced decision-making in complex multi-agent environments. Furthermore, our findings contribute to the broader field of machine learning by offering insights into managing mixed-motivation scenarios. Consequently, our work paves the way for future research and innovation, driving progress in multi-agent learning and its applications across various domains.
Appendix A Proof of Corollary 4.3
@@
We then extend the definition of infinitesimal alignment (Balduzzi et al., 2018) to our proposed AgA gradient. Given a third reference vector ${\bm{w}}$, the infinitesimal alignment for AgA gradient $\tilde{{\bm{\xi}}}$ with ${\bm{w}}$ is defined as
@@
Lemma A.1 (Sign of $align$.).
@@
Lemma A.1 empowers us to calculate the sign of infinitesimal alignment relying on components that are readily computable.
@@
1) Case 1: In a neighborhood of a stable fixed point. If we are in a neighborhood of a stable fixed point then $\langle{\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c}\rangle\geq 0$. This indicates that both vectors ${\bm{\xi}}_{c}$ and $\nabla{\mathcal{H}}_{c}$ point in the same direction, i.e., $\theta({\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c})\leq\pi/2$. Referring to Lemma A.1, the sign of $align(\tilde{{\bm{\xi}}},\nabla{\mathcal{H}}_{c})$ is the same as the sign of $\langle{\bm{\xi}},\nabla{\mathcal{H}}_{c}\rangle+\|\nabla{\mathcal{H}}_{c}\|^{2}$ if $\langle{\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c}\rangle\geq 0$. Consequently, we have $sign(\lambda)=sign(\langle{\bm{\xi}},\nabla{\mathcal{H}}_{c}\rangle+\|\nabla{\mathcal{H}}_{c}\|^{2})=sign(align)$ matching the claim of our corollary.
@@
Therefore, we prove that if we let the sign of $\lambda$ follows the condition $\lambda\cdot\langle{\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c}\rangle(\langle{\bm{\xi}},\nabla{\mathcal{H}}_{c}\rangle+\|\nabla{\mathcal{H}}_{c}\|^{2})\geq 0$, the optimization process using the AgA method exhibit following attributes. In areas close to fixed points, 1) if the point is stable, the AgA gradient will be pulled toward this point, which means that $\theta(\tilde{{\bm{\xi}}},\nabla{\mathcal{H}}_{c})\leq\theta({\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c})$; 2) if the point represents unstable equilibria, the AgA gradient will be pushed out of the point, indicating that $\theta(\tilde{{\bm{\xi}}},\nabla{\mathcal{H}}_{c})\geq\theta({\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c})$. An illustrative example of the corollary is provided in Fig. 2.
Appendix B Additional Experiment Results

### v2

@@ 5.1 Two-Player Public Goods Game
A two-player public goods matrix game ${\mathcal{G}}$ is widely utilized to study cooperation in social dilemmas. The game involves players $\{1,2\}$ with parameters $\{{\bm{w}}_{1},{\bm{w}}_{2}\}$ and payoffs $\{p_{1},p_{2}\}$. The social welfare, denoted as $SW=p_{1}+p_{2}$. Each player, i.e., $i\in\{1,2\}$, contributes an amount $a_{i}$ within a budgeted range $[0,b]$, and the host evenly distributes these contributions as $\frac{c}{2}(a_{1}+a_{2})$, where $1<c\leq 2$. Consequently, each player’s payoff $p_{i}(a_{1},a_{2})$ is estimated as $b-a_{i}+\frac{c}{2}(a_{1}+a_{2})$. In our experiments, we set the budget $b$ to $1$ and weight $c$ to $1.5$, with the social optimum of the game at $(1,1)$.
Table 1: The comparison of the average individual rewards (denoted as $r_{1},r_{2}$), social welfare (denoted as $SW$) on two- player public goods game. We show the mean of rewards and 95% confidence interval utilizing 50 random runs.
Metrics Simul-Ind CGA SGA SVO Simul-Co SL AgA
$\mathbf{r_{1}}$ 1.133 $\pm$ 0.063 1.156 $\pm$ 0.060 1.175 $\pm$ 0.062 1.104 $\pm$ 0.054 1.433 $\pm$ 0.056 1.314 $\pm$ 0.062 1.443$\pm$ 0.042
$\mathbf{r_{2}}$ 1.184 $\pm$ 0.065 1.150 $\pm$ 0.057 1.137 $\pm$ 0.063 1.060 $\pm$ 0.051 1.381 $\pm$ 0.065 1.371 $\pm$ 0.057 1.459 $\pm$ 0.041
$\mathbf{SW}$ 2.316$\pm$ 0.039 2.306 $\pm$ 0.039 2.312$\pm$ 0.044 2.164 $\pm$ 0.026 2.814 $\pm$ 0.033 2.684 $\pm$ 0.049 2.903$\pm$ 0.023
Results. Table 1 presents a comparison of individual rewards and the collective outcome, derived from 50 random trials, each limited to 100 steps. The rows $r_{1},r_{2}$ and $SW$ represent the rewards accrued by each individual player and the collective group, respectively. The social optimum of the game occurs when all players allocate their entire budget, achieving the highest possible social welfare score of $3$. Among all algorithms, AgA achieves the closest social welfare score to this optimum, with a value of 2.90. Furthermore, AgA exhibits the greatest degree of fairness, with two players receiving nearly identical rewards, in contrast to the baseline algorithms. Furthermore, Fig. 4 in Appendix D illustrates the action distributions of these algorithms, showing that AgA actions are more tightly clustered around the social optimum actions $(1,1)$. This further indicates that AgA approaches the social optimum more effectively.
(a) Cleanup
(b) Harvest
(c) Selfish-MMM2
(d) Cleanup
(e) Harvest
(f) Selfish-MMM2
(g) Cleanup
(h) Harvest
(i) Selfish-MMM2
Figure 3: The first row displays results comparing different values of the alignment parameter $\lambda$ across three environments: Cleanup and Harvest (measurement of social welfare, SW) and Selfish-MMM2 (focusing on win rate). The second row examines the performance differences between the proposed AgA and AgA without sign alignment (AgA-Sign) on the three testbeds. The bold lines indicate the mean social welfare calculated in three seeds, while the surrounding shaded areas represent the 95% confidence interval. The third row compares the AgA method with baseline approaches on these testbeds. Each bar represents the mean collective results of each method and the error bars indicate the 95% confidence interval.
@@ 5.2 Sequential Social Dilemma: Cleanup and Harvest
We then verify the AgA algorithm in two widely used sequential social dilemma games with 5 homogeneous agents: Harvest and Cleanup (Hughes et al., 2018). In the Cleanup scenario, agents predominantly gather rewards by harvesting apples in an orchard, where the yield is affected by the river’s pollution levels. Neglecting rising pollution levels ultimately ceases apple production, thus setting up a trade-off between individual gains and communal welfare. In contrast, the Harvest scenario compensates agents for apple collection, with the regeneration of apples being ideally dependent on the proximity to other apples. Here, the communal challenge manifests itself as over-harvesting apples, which diminishes their regrowth rate and, consequently, the aggregate rewards for the group. Our research assessed AgA against various benchmarks, including Simul-Ind, Simul-Co, CGA, SVO, and SL. Simul-Ind was implemented using the IPPO algorithm (de Witt et al., 2020), while the remaining benchmarks employed the common parameters of the IPPO framework. The formula for calculating the collective loss for CGA and AgAs, which aims to improve both the performance of the group and the equitable distribution between agents, is expressed as $\sum_{i}\left(r_{i}-\alpha(1-\arctan\left(\frac{\sum_{j,j\neq i}r_{j}}{r_{i}}\right)\right)$, where $\alpha$ is a constant. Further details on the experimental setups and the algorithms used are provided in Appendix E.1.
Varying $\lambda$ values. Fig. 3(a) and Fig. 3(b) explore the social welfare outcomes from AgA under three distinct $\lambda$ values: 0.1, 1, 100, and 1000. Among these, $\lambda=100$ demonstrates superior performance in both the Cleanup and Harvest scenarios. Therefore, in the subsequent experiments involving sign alignment and comparing primary results with baseline models, we will employ $\lambda=100$ as the default parameter.
The effectiveness of sign alignment. In the second row of Fig. 3, Fig. 3(d) and Fig. 3(e) illustrate the social welfare comparisons between AgA and AgA without sign alignment (AgA-Sign) during their training phases. Referencing Corollary 4.3, near a fixed point, sign alignment helps direct the gradient towards stable fixed points and away from unstable ones. During the latter half of the training sessions shown in Fig. 3(d) and Fig. 3(e), it is observed that the social welfare ($SW$) for AgA-Sign does not improve to the same extent as it does for AgA, highlighting the effectiveness of sign alignment.
Baseline Comparison. As illustrated in Fig. 3(g) and Fig. 3(h), AgA demonstrates a notable improvement in social welfare over Simul-Ind, Simul-Co, CGA, SVO, and SL. Particularly, within the Cleanup scenario, AgA recorded an average social welfare of 105.15, marking approximately a 56% increase over the runner-up method SL. In the Harvest scenario, AgA shows an average improvement of 33.55 in social welfare over the next best method SVO.
5.3 Selfish-MMM2
While sequential social dilemma games are frequently used to study mixed-motive problems, they are relatively simplistic compared to other experimental environments in MARL, particularly when considering real-world applications. To address this gap, we introduce a sophisticated mixed-motive testbed called Selfish-MMM2, which is adapted from the MMM2 map in the StarCraft II game (Samvelyan et al., 2019). Unlike the standard SMAC framework, which employs a shared reward structure, our methodology assigns individual objectives to each agent, thereby enhancing the complexity of mixed motives. Specifically, we have redesigned the reward systems so that each agent’s personal gain is directly tied to the damage they inflict on adversaries, diverging from traditional approaches where collective damage is evenly distributed among all agents. Additionally, penalties for agent elimination emphasize the importance of self-preservation and highlight agents’ tendencies towards self-interested behavior. In comparison to widely utilized multi-agent environments such as Cleanup (Hughes et al., 2018), Harvest (Hughes et al., 2018), and the 10-player Prisoner’s Dilemma (PD) for large-scale evaluations (Gemp et al., 2020), Selfish-MMM2 is distinguished by two notable features: 1) Large-Scale Agents: Selfish-MMM2 manages the interactions of 10 heterogeneous agents, categorized into three types: seven Marines, two Marauders, and one Medivac, confronting 11 enemy units. This configuration is in stark contrast to Cleanup and Harvest, which each utilize five homogeneous agents. 2) Extensive Action Space: The action space in Selfish-MMM2, sized at $18^{10}$, is vastly larger than that of Cleanup and Harvest, which are limited to $9^{5}$ and $8^{5}$, respectively. It also surpasses the action space of the 10-player PD, which is $2^{10}$.
In evaluating the Selfish-MMM2 environment, we use the average win rate against the built-in AI bot in SMAC to measure the collective performance. The experimental setup employs the IPPO algorithm (de Witt et al., 2020) for the Simul-Ind algorithm, while the MAPPO (Yu et al., 2022) is utilized for the Simul-Co, SVO, CGA, SL and AgA algorithms. More details on the Selfish-MMM2 environment and implementation specifics are available in the Appendix E.2.
Varying $\lambda$ Values and the effectiveness of Sign Alignment: Fig. 3(c) presents the average win rates in various settings of the align parameter $\lambda$ in the AgA scenario. The findings indicate that increased $\lambda$ values are associated with faster convergence in the training period. Among these, a $\lambda$ value of 100 demonstrates a higher overall effectiveness compared to other values tested. Based on these findings, we have set a default value $\lambda$ of 100 for subsequent experiments. Furthermore, Fig. 3(f) in the second row illustrates a comparable conclusion in sequential social dilemma games. During the initial training phase, the convergence rates of AgA and AgA-Sign are nearly identical. However, as training progresses, AgA exhibits convergence at points of superior performance, highlighting the critical role of sign alignment in achieving convergence near fixed points.
Baseline Comparison. As depicted in Fig. 3(i), the bar chart displays a comparison of the average win rates along with 95% confidence intervals for AgA versus other baseline approaches. Our AgA approach attains the highest team performance, securing a 92.64% win rate with the built-in bot, which is 23.61 percent more than the next best win rate achieved by the SL method.
@@ 6 Conclusion
In this paper, we propose AgA, an optimization method specifically designed to align individual and collective objectives through gradient adjustments in mixed-motive cooperation scenarios. To achieve this, we first model the mixed-motive game as a differentiable game, offering a powerful tool for analyzing learning dynamics at both individual and collective levels. Furthermore, we theoretically demonstrate that AgA can effectively attract gradients to stable fixed points and support our claims with empirical evidence. To evaluate the effectiveness of AgA in complex and large-scale scenarios, we introduce a new mixed-motive environment called Selfish-MMM2, which features heterogeneous large-scale agents, a larger action space, and increased task complexity. Comprehensive experiments, ranging from simple public goods games to Harvest, Cleanup, and Selfish-MMM2, show AgA’s superior performance, consistently outperforming existing baselines across multiple evaluation metrics. These results validate our theoretical claims and highlight the effectiveness of AgA in achieving cooperative behavior in multi-agent systems.
Limitations. Despite conducting a series of experiments to verify the proposed AgA algorithm, including using a new large-scale mixed-motive cooperation testbed, these environments remain somewhat removed from real-world applications. Moreover, while we do not explicitly derive the Hessian matrix in AgA, the algorithm introduces an additional computational complexity of $\mathcal{O}(n)$ to compute the Hessian-vector product. Although we strive to align the objectives of both individuals and the collective, our primary focus in this paper is on converging towards stable fixed points of the collective objective. For future research, we plan to delve deeper into the interaction of individual objectives in mixed-motive games, with an increased emphasis on understanding their dynamics in real-world mixed-motive cooperation scenarios.
Appendix A Broader Impacts
@@
This paper presents a novel approach in the field of multi-agent learning, aimed at enhancing collaboration among agents with mixed motivations. The proposed framework addresses a fundamental challenge in machine learning: enabling agents with diverse objectives to cooperate effectively to achieve collective goals. The work described in this paper holds significant potential for various societal applications. By facilitating effective cooperation among agents with different motivations, our approach can be applied to domains such as video games, autonomous systems, smart cities, and decentralized resource allocation. These advancements can lead to numerous societal benefits, including improved efficiency in transportation systems, optimized resource allocation in energy grids, and enhanced decision-making in complex multi-agent environments. Furthermore, our findings contribute to the broader field of machine learning by offering insights into managing mixed-motivation scenarios. This work paves the way for future research and innovation, driving progress in multi-agent learning and its applications across various domains. However, while our method was initially developed to address fundamental challenges in decision-making within machine learning contexts, it is conceivable that our proposed objective alignment techniques could be repurposed for harmful purposes, such as attacking existing cooperative systems.
Appendix B Proof of Corollary 4.3
@@
We then extend the definition of infinitesimal alignment [Balduzzi et al., 2018] to our proposed AgA gradient. Given a third reference vector ${\bm{w}}$, the infinitesimal alignment for AgA gradient $\tilde{{\bm{\xi}}}$ with ${\bm{w}}$ is defined as
@@
Lemma B.1 (Sign of $align$.).
@@
Lemma B.1 empowers us to calculate the sign of infinitesimal alignment relying on components that are readily computable.
@@
1) Case 1: In a neighborhood of a stable fixed point. If we are in a neighborhood of a stable fixed point then $\langle{\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c}\rangle\geq 0$. This indicates that both vectors ${\bm{\xi}}_{c}$ and $\nabla{\mathcal{H}}_{c}$ point in the same direction, i.e., $\theta({\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c})\leq\pi/2$. Referring to Lemma B.1, the sign of $align(\tilde{{\bm{\xi}}},\nabla{\mathcal{H}}_{c})$ is the same as the sign of $\langle{\bm{\xi}},\nabla{\mathcal{H}}_{c}\rangle+\|\nabla{\mathcal{H}}_{c}\|^{2}$ if $\langle{\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c}\rangle\geq 0$. Consequently, we have $sign(\lambda)=sign(\langle{\bm{\xi}},\nabla{\mathcal{H}}_{c}\rangle+\|\nabla{\mathcal{H}}_{c}\|^{2})=sign(align)$ matching the claim of our corollary.
@@
Therefore, we prove that if we let the sign of $\lambda$ follows the condition $\lambda\cdot\langle{\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c}\rangle(\langle{\bm{\xi}},\nabla{\mathcal{H}}_{c}\rangle+\|\nabla{\mathcal{H}}_{c}\|^{2})\geq 0$, the optimization process using the AgA method exhibit following attributes. In areas close to fixed points, 1) if the point is stable, the AgA gradient will be pulled toward this point, which means that $\theta(\tilde{{\bm{\xi}}},\nabla{\mathcal{H}}_{c})\leq\theta({\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c})$; 2) if the point represents unstable equilibria, the AgA gradient will be pushed out of the point, indicating that $\theta(\tilde{{\bm{\xi}}},\nabla{\mathcal{H}}_{c})\geq\theta({\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c})$. An illustrative example of the corollary is provided in Fig. 2(a).
Appendix C Implementation of MARL Algorithms with AgA
The concrete implementation of altruistic gradient adjustment is elaborated in Algorithm 1. AgA can be easily incorporated into any centralized training decentralized execution (CTDE) framework of MARL. In a typical CTDE framework, each agent, denoted as $i$, strives to learn a policy that employs local observations to prompt a distribution over personal actions. The unique characteristic during the centralized learning phase is that agents gain supplementary information from their peers that is inaccessible during the execution stage. Furthermore, a centralized critic often operates in tandem in CTDE settings, leveraging shared information and evaluating individual agent policies’ potency. To incorporate our proposed AgA algorithm into CTDE frameworks, supplementary reward exchanges between agents are required, offering the essential inputs for Algorithm 1. The individual losses are calculated from individual rewards conforming to established MARL algorithm framework, such as IPPO [de Witt et al., 2020] or MAPPO [Yu et al., 2022]. Here, the collective loss $\ell_{c}$ is calculated according to individual rewards; for example, it could be a sum of individual rewards or an elementary transformation thereof. Furthermore, the AgA algorithm requires a magnitude value of the alignment parameter $\lambda$, which should be a positive entity. The sign of $\lambda$ is then calculated in line 5 of Algorithm 1 and is based on each gradient component calculated in lines 2-4. Subsequently, the algorithm produces the adjusted gradient $\tilde{{\bm{\xi}}}$. $\tilde{{\bm{\xi}}}$ is then distributed across each corresponding parameter, and optimization is carried out using a suitable optimizer such as Adam optimizer [Kingma and Ba, 2015].
Algorithm 1 Altruistic Gradient Adjustment (AgA)
3:  ${\bm{\xi}}_{c}\leftarrow[grad(\ell_{c},{\bm{w}}_{i})\ for\ {\bm{w}}_{i}\in{\bm{w}}]$
Appendix D Additional Experiment Results